# Quasi-steady emission from repeating fast radio bursts can be explained by magnetar wind nebula

### v1

@@ 6 Summary & Discussions
Our optimal parameter values for $B_{\rm dip}$, $P_{i}$, $t_{\rm age}$, $M_{\rm ej}$ and $E_{\rm sn}$ are consistent with those derived from previous studies (see e.g., Kashiyama et al. 2016; Yang & Dai 2019; Zhao & Wang 2021). For a pair-dominated wind, Yang & Dai (2019) proposed that persistent radio counterparts to FRBs could arise from MWN powered by a magnetar without surrounding SN ejecta. However, their model could not explain the time evolution of source DM and free-free absorption in the SN ejecta. Recently, Chen & Tong (2024) used a one-zone PWN model to explain the observed quasi-steady emission from R1 and R2. Both these works, however, assumed a constant energy injection rate into a freely expanding nebula. In contrast, Margalit & Metzger (2018) and Zhao & Wang (2021) modelled the energy injection into MWN using a more realistic time-dependent function similar to our work. The predicted $t_{\rm age}\sim 10-40\,{\rm yr}$ and $E_{\rm sn}\sim 10^{50}-10^{51}\,{\rm erg}$ for FRB 121102 from Margalit & Metzger (2018) agrees well with our estimates. Similarly, Zhao & Wang (2021) derived $B_{\rm dip}\approx 0.1\,B_{\rm int}\sim(2-4)\times 10^{15}\,{\rm G}$ and $t_{\rm age}\sim 14-22\,{\rm yr}$ from the radio SEDs of R1 and R2. In both these works, the dominant energy injection channel into the nebula was intermittent magnetar flares and a constant nebula/ejecta velocity was assumed. In our analysis, we improved upon these works by first deriving constraints on the NS and ejecta parameters using observed properties (DM, energy, PRS size, radio signal non-attenuation), to then self-consistently compute the synchrotron flux generated by energized electrons considering both NS rotational and magnetic energy injection scenarios, for nebula/ejecta velocity that is driven by the time-dependent energy injection rate into the magnetized nebula.
@@
If repeating FRBs were to arise from host galaxies that preferentially harbor SLSNe, their rates would be smaller $\sim 0.01\%$ of CCSNe (Quimby et al., 2013). Conversely, if these sources were associated with either ultra-stripped SNe or AICs, their rates should be somewhat higher $\sim 0.1-1\%$ of CCSNe (Ruiter et al., 2009; Tauris et al., 2013). Once the number density of repeating FRBs is well constrained from observations, it would facilitate the distinction between these possibilities. Future searches for radio counterparts of pulsar-driven SN candidates, including SLSNe, should give us crucial information on their relationship with FRBs. This is especially true for luminous transients with rapid rotation $P_{i}\sim 1-30\,{\rm ms}$, that are powered by the spin-down luminosity of the NS rather than its magnetic energy. Such nebular emission would be non-thermal, and therefore, easily distinguishable from Galactic thermal sources.
@@
Recently, Xing et al. (2024) reported the detection of a GeV gamma-ray flare lasting $\sim 15\,{\rm sec}$ associated with the hyper-active repeating FRB 20240114A. Our MWN model naturally predicts the existence of such energetic gamma-ray flares in the forward shock region which can result from magnetic energy dissipation within the nebula (see also Murase et al. 2016). Furthermore, hard X-ray emission can also serve as another promising signal for the parameters that account for both SLSNe and FRBs (Metzger & Piro, 2014; Kashiyama et al., 2016). A fraction of X-rays can be absorbed by SN ejecta to be re-emitted as thermal emission leading to an optical/UV transient on timescales of days. However, the optical depth of pairs in the nebula can suppress the thermalization efficiency of $L_{\rm sd}$. For the cases where nebular X-rays are sufficiently energetic to reionize the ejecta, non-thermal X-rays can escape the ejecta unattenuated with a peak luminosity and timescale that is similar to optical radiation.
@@ Acknowledgements
We thank Ke Fang and Pawan Kumar for useful discussions and comments. M.B. acknowledges support from the Eberly Research Fellowship at the Pennsylvania State University and the Simons Collaboration on Extreme Electrodynamics of Compact Sources (SCEECS) Postdoctoral Fellowship at the Wisconsin IceCube Particle Astrophysics Center (WIPAC), University of Wisconsin-Madison. The work of K.M. is supported by the NSF Grant No. AST-2108466, No. AST-2108467, and AST-2308021, and KAKENHI No. 20H05852.

### v2

@@ 6 Discussion
Our optimal parameter values ($B_{\rm dip}$, $P_{i}$, $t_{\rm age}$, $M_{\rm ej}$, $E_{\rm sn}$) are consistent with previous studies (e.g., Kashiyama et al. 2016; Yang and Dai 2019; Zhao and Wang 2021). Some works proposed models without SN ejecta (Yang and Dai, 2019) or used simplified one-zone models with constant energy injection (Chen and Tong, 2024), which face challenges explaining DM evolution or absorption. Others modelled time-dependent injection, often focusing on magnetar flares and assuming constant nebula/ejecta velocity (Margalit and Metzger, 2018; Zhao and Wang, 2021). Our analysis improves upon these studies by first deriving constraints on NS and ejecta parameters from observations (DM, energy, PRS size, signal attenuation), then self-consistently computing the synchrotron flux for both NS rotational and magnetic energy injection scenarios, using nebula/ejecta velocities driven by the time-dependent energy injection rate into the MWN.
We note that the present framework focuses primarily on radiative evolution of the magnetar wind nebula, with the dynamical and acceleration components treated in a simplified, phenomenological manner. A fully self-consistent model would require simultaneous treatment of the wind–ejecta interaction and particle acceleration for arbitrary magnetization, which is beyond the scope of this work. Our parameterization in terms of $\epsilon_{e}$, $\epsilon_{B}$ and $\gamma_{\rm e,min}$ effectively assumes efficient conversion of upstream Poynting flux to downstream particle energy. The inferred magnetar parameters should therefore be interpreted as conditional upon these assumptions. While quantitative constraints may vary for different dissipation efficiencies or acceleration prescriptions, the qualitative conclusion that young magnetars embedded in synchrotron-emitting nebulae can account for the observed luminosities and spectra of PRSs remains robust.
@@
If repeating FRBs arise from host galaxies that preferentially harbor SLSNe, their rates would be low ($\sim 0.01\%$ of CCSNe, Quimby et al. 2013). Conversely, if these were associated with USSN or AICs, rates might be higher ($\sim 0.1-1\%$ of CCSNe, Ruiter et al. 2009; Tauris et al. 2013). Constraining the number density of repeating FRBs will help distinguish these possibilities. Future searches for radio counterparts of pulsar-driven SN candidates, including SLSNe, especially those powered by rapid NS rotation ($P_{i}\sim 1-30\,{\rm ms}$), will give crucial information on their relationship with FRBs. Such non-thermal nebular emission should be distinguishable from Galactic thermal sources.
@@
Recently, Xing et al. (2024) reported a GeV gamma-ray flare associated with the hyper-active repeating FRB 20240114A. While our MWN model allows for such emission in the forward shock region (Murase et al., 2016), reproducing the observed gamma-ray flux requires more efficient magnetic energy dissipation than assumed here. Furthermore, hard X-ray emission can also serve as another promising signal for parameters accounting for both SLSNe and FRBs (Metzger and Piro, 2014; Kashiyama et al., 2016). While some X-rays might be reprocessed by SN ejecta into thermal optical/UV transient lasting over days, the optical depth of pairs in the nebula could suppress thermalization. If nebular X-rays are energetic enough to reionize the ejecta, non-thermal X-rays could escape unattenuated with a peak luminosity and timescale similar to optical radiation.
7 Summary
This study investigated the origin of PRSs associated with three localized repeating FRBs: 20121102A, 20190520B, and 20201124A. Recently, two more confirmed FRB-PRS associations have been reported for FRBs 20240114A and 20190417A. Motivated by the potential for young and rapidly rotating magnetars powering wind nebulae (MWN) to explain various energetic transients such as superluminous supernovae (SLSNe), we explored the scenario where the observed persistent radio emission arises due to quasi-steady synchrotron radiation from an MWN expanding within the SN ejecta that birthed the magnetar.
We employed a detailed theoretical model, numerically solving kinetic equations for particles and photons within the MWN to predict radio SEDs and light curves, considering emission/absorption processes and electromagnetic cascades. Both NS spindown and magnetar flares were treated as possible energy sources. Our model is tested against two distinct SN progenitor environments – an ultra-stripped SN with low ejecta mass and a conventional core-collapse SN with high ejecta mass – while constraining key NS parameters ($P_{i},B_{\rm dip},t_{\rm age}$) using physical limits derived from energy, dispersion measure, radio attenuation, and nebula size.
Our results indicated that the NS rotational energy is generally sufficient to power observed PRS luminosities. Comparing model predictions with observations, we found the properties of FRBs 20121102A and 20190520B (higher flux at lower frequencies, harder spectra) are better matched by a magnetar aged $\sim 20$ yr within a USSN environment. Conversely, the distinct low-frequency turnover in FRB 20201124A’s spectrum points towards significant self-absorption, favoring a younger magnetar ($\sim 10-12.5$ yr) residing in the denser ejecta of a CCSN. These findings are consistent with minimum age estimates of $\sim 7-10$ yr derived from radio transparency constraints, lending support to the MWN model as a viable explanation for FRB persistent radio counterparts and potentially linking some FRBs to SN progenitor channels.
@@ Acknowledgements
We thank Ke Fang, Pawan Kumar and Navin Sridhar for useful discussions and comments. We also thank the anonymous reviewer for their helpful comments and suggestions that improved the quality of this manuscript. M.B. acknowledges support from the Eberly Research Fellowship at the Pennsylvania State University and the Simons Collaboration on Extreme Electrodynamics of Compact Sources (SCEECS) Postdoctoral Fellowship at the Wisconsin IceCube Particle Astrophysics Center (WIPAC), University of Wisconsin-Madison. The work of K.M. is supported by the NSF Grant No. AST-2108466, No. AST-2108467, and AST-2308021, and KAKENHI No. 20H05852. The work of K.K. is supported by JSPS KAKENHI Grant Nos. JP24K00668, JP23H04899, JP22H00130.